OpenStudy (anonymous):

Please help! I was trying to integrate this then i got stuck dr/dt= 2r(1-r). So far i have: log(r)-log(1-r)+c=2t. I know I'm supposed to use e^log but i'm confused as to how to do it

OpenStudy (jamesj):

hence \[\ln\left(\frac{r}{1-r}\right) = 2t - c\]

OpenStudy (jamesj):

Now take the exp of both sides and remember that exp(ln(a)) = a

OpenStudy (jamesj):

I.e., \[ {r \over 1 - r} = Ae^{2t} \]

OpenStudy (jamesj):

where A = e^-c Now solve for r

OpenStudy (anonymous):

Thank youuuu!